# This is the extended version of the conference version at FoSSaCS 2022 [BFPZ22] [a] [a] [a]

### v1

@@ 6.1.1. Checking causality and existence of reachability PR causes
$\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\lozenge\mathsf{Cause}\wedge\mathsf{rEff})$ $=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}}(\lozenge\mathsf{eff}_{\mathsf{cov}})=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}}(\lozenge\mathsf{Cause}_{\mathcal{A}_{\mathsf{rEff}}}\wedge\lozenge E).$ (27)
Hence, $\mathsf{Cause}^{\prime}$ is a GPR cause for $\{\mathsf{eff}_{\mathsf{cov}},\mathsf{eff}_{\mathsf{unc}}\}$ in $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ if and only if $\mathsf{Cause}$ is a reachability GPR cause for $\mathsf{rEff}$ in $\mathcal{M}$.
@@
This Lemma 42 shows that reachability SPR causality shares similarities with GPR causality. As an immediate consequence we can check conditions (rG) and (rS) in $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ by using the already established algorithms for GPR causes. This results in the following complexity upper bounds.
@@ 6.1.2. Computing quality measures of reachability PR causes
where the values for $\mathfrak{T}$ are defined in $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ as in Section 5.1.
@@ 6.1.3. Finding quality optimal reachability PR causes
Intuitively, this states that it is never already certain that the execution will belong to $\mathsf{rEff}$ when the cause is reached. A variation of this criteria has been proposed also in [BDF+21].
@@
Now, we want to find recall- and coverage ratio-optimal reachability SPR causes. As in the state-based setting, we define the set $\mathcal{C}$ of all possible singleton reachability SPR causes for $\mathsf{rEff}$ that in addition satisfy (TempPrio) as explained in Remark 47. By Corollary 43, we can check whether a state $c\in S$ is a singleton reachability SPR cause in $\mathrm{coNP}$; whether there is at least one state with $c$ in the first component in $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ such that $\mathsf{noeff}_{\mathsf{fp}}$ is reachable from this state is checkable in polynomial time. As before, the canonical cause $\mathsf{CanCause}$ is now the set of states $c\in\mathcal{C}$ for which there is a scheduler $\mathfrak{S}$ with $\mathrm{Pr}^{\mathfrak{S}}(\neg\mathcal{C}\,\mathrm{U}\,c)$.
@@ 6.2.1. Checking co-safety causality
Let now $\mathcal{M}$ be an MDP, $\mathsf{rEff}$ an effect given by the DRA $\mathcal{A}_{\mathsf{rEff}}$ and $\mathsf{rCause}$ a cause candidate given by a DFA $\mathcal{A}_{\mathsf{rCause}}$ as above. So, in particular, $\mathcal{A}_{\mathsf{rCause}}$ accepts exactly the minimal good prefixes for $\mathsf{rCause}$. For the special case of Markov chains the check can be done in polynomial time analogously to reachability PR causes by computing the corresponding conditional probabilities. We can provide a model transformation of $\mathcal{M}$ using both automata such that the resulting MDP has no end components and the effect is a reachability property again similar to Notation 6.1.1.
@@
Therefore, the desired reduction as in Lemma 50 does not work for $\mathcal{M}$. Note that the violation of the condition (coS) is only possible in this example if the scheduler behaves differently depending on how state $c$ is reached. This different behavior, however, does not have anything to do with the effect, and potentially different residual properties that have to be satisfied to achieve the effect; in the example, the effect is just a reachability property. Furthermore, we want to emphasize that the concrete probabilities of the individual paths leading to $c$ are important for the violation. In general, this imposes a major challenge for checking the condition (coS), for which we do not know a solution. Similar problems arise when trying to check the existence of a co-safety SPR cause. A witness might be just one individual path, potentially only together with a scheduler that realizes this path with very low probability.
@@ 6.2.3. Finding optimal co-safety PR causes
Figure 12. A Markov chain $\mathcal{M}$ in which the regular effect $\mathsf{rEff}=\lozenge e$ has no recall-optimal regular PR-cause.
@@ 7. Conclusion
The approach of probability-raising within this work is in line with the classical notion in literature that uses a strict inequality in the PR condition. As a consequence causes might not exist (see Example 2). However, relaxing the PR condition by only requiring a non-strict inequality would apparently be a minor change that broadens the choice of causes. Indeed, the proposed algorithms for checking the SPR and GPR condition for reachability effects (Section 4) can easily be modified for the relaxed definition. As the algorithms of both extended settings discussed in Section 6 stem mainly from a reduction to reachability effects this also holds for reachability and co-safety causes of regular effects. However, a non-strict inequality in the PR condition would lead to a questionable notion of causality, as e.g. $\{\mathsf{init}\}$ would always be a recall- and ratio-optimal cause. Thus, other side constraints are be needed in order to make use of the relaxed PR condition. E.g., requiring the non-strict inequality for all schedulers that reach a cause with positive probability and also requiring the existence of a witnessing scheduler for the PR condition with strict inequality might be a useful alternative definition which agrees with Def. 3 for Markov chains.
@@ Acknowledgments
We would like to thank Florian Funke, Simon Jantsch and Clemens Dubslaff for their helpful comments and feedback on the topic of causality in MDPs. We especially thank Simon Jantsch for pointing out improvements of prior results (wrt. [BFPZ22]).

### v2

@@ 6.1.1. Checking causality and existence of reachability PR causes
$\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\lozenge\mathsf{Cause}\wedge\mathsf{rEff})$ $=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}}(\lozenge\mathsf{eff}_{\mathsf{cov}})=\mathrm{Pr}^{\mathfrak{T}}_{\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}}(\lozenge\mathsf{Cause}_{\mathcal{A}_{\mathsf{rEff}}}\wedge\lozenge E),$ (27)
where $E=\{\mathsf{eff}_{\mathsf{cov}},\mathsf{eff}_{\mathsf{unc}}\}$. Hence, $\mathsf{Cause}_{\mathcal{A}_{\mathsf{rEff}}}$ is a GPR cause for $\{\mathsf{eff}_{\mathsf{cov}},\mathsf{eff}_{\mathsf{unc}}\}$ in $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ if and only if $\mathsf{Cause}$ is a reachability GPR cause for $\mathsf{rEff}$ in $\mathcal{M}$.
@@
This Lemma 42 shows that reachability SPR causality shares similarities with GPR causality. Our algorithmic results for reachability PR causes stem from the reduction provided by Notation 6.1.1 and Lemma 42. As an immediate consequence we can check conditions (rG) and (rS) in $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ by using the already established algorithms for GPR causes. This results in the following complexity upper bounds.
@@ 6.1.2. Computing quality measures of reachability PR causes
where the values for $\mathfrak{T}$ are defined in $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ as for the state-based setting (cf. Section 5.1).
@@ 6.1.3. Finding quality optimal reachability PR causes
Intuitively, this states that it is never already certain that the execution will belong to $\mathsf{rEff}$ when the cause is reached. A variation of this criterium has been proposed also in [BDF+21].
@@
Now, we want to find recall- and coverage ratio-optimal reachability SPR causes. As in the state-based setting, we define the set $\mathcal{C}$ of all possible singleton reachability SPR causes for $\mathsf{rEff}$ that in addition satisfy (TempPrio) as explained in Remark 47. By Corollary 43, we can check whether a state $c\in S$ is a singleton reachability SPR cause in coNP; whether there is at least one state with $c$ in the first component in $\mathcal{M}_{[\mathsf{rEff},\mathsf{Cause}]}$ such that $\mathsf{noeff}_{\mathsf{fp}}$ is reachable from this state is checkable in polynomial time. Thus, we can again define the set of singleton reachability SPR causes $\mathcal{C}=\{s\in S\mid s\text{ is reachability SPR cause satisfying }\mathrm{Pr}^{\min}_{\mathcal{M}}(\mathsf{rEff}\mid\neg\mathsf{Cause}\,\mathrm{U}\,c)<1\}$. As before, the canonical cause $\mathsf{CanCause}$ is now the set of states $c\in\mathcal{C}$ for which there is a scheduler $\mathfrak{S}$ with $\mathrm{Pr}^{\mathfrak{S}}(\neg\mathcal{C}\,\mathrm{U}\,c)$.
@@ 6.2.1. Checking co-safety causality
Let now $\mathcal{M}$ be an MDP, $\mathsf{rEff}$ an effect given by the DRA $\mathcal{A}_{\mathsf{rEff}}$ and $\mathsf{rCause}$ a cause candidate given by a DFA $\mathcal{A}_{\mathsf{rCause}}$ as above. So, in particular, $\mathcal{A}_{\mathsf{rCause}}$ accepts exactly the minimal good prefixes for $\mathsf{rCause}$. We now want to check, whether $\mathsf{rCause}$ is a co-safety SPR cause (resp. co-safety GPR cause). For the special case of Markov chains the check can be done in polynomial time analogously to reachability PR causes by computing the corresponding conditional probabilities. We can provide a model transformation of $\mathcal{M}$ using both automata such that the resulting MDP has no end components and the effect is a reachability property again similar to Notation 6.1.1.
@@
[…]
@@ 6.2.3. Finding optimal co-safety PR causes
Figure 18. Markov chain $\mathcal{M}$ where the regular effect $\lozenge e$ has no recall-optimal regular PR-cause.
@@ 7. Conclusion
The approach of probability-raising within this work is in line with the classical notion in literature that uses a strict inequality in the PR condition. As a consequence causes might not exist (see Example 2). However, relaxing the PR condition by only requiring a non-strict inequality would apparently be a minor change that broadens the choice of causes. Indeed, the proposed algorithms for checking the SPR and GPR condition for reachability effects (Section 4) can easily be modified for the relaxed definition. As the algorithms of both extended settings discussed in Section 6 stem mainly from a reduction to reachability effects this also holds for reachability and co-safety causes of regular effects. However, a non-strict inequality in the PR condition would lead to a questionable notion of causality, as e.g. $\{\mathsf{init}\}$ would always be a recall- and ratio-optimal cause. Thus, other side constraints are needed in order to make use of the relaxed PR condition. E.g., requiring the non-strict inequality for all schedulers that reach a cause with positive probability and also requiring the existence of a witnessing scheduler for the PR condition with strict inequality might be a useful alternative definition which agrees with Def. 3 for Markov chains.
@@ Acknowledgments
We would like to thank Florian Funke and Clemens Dubslaff for their helpful comments and feedback on the topic of causality in MDPs. We especially thank Simon Jantsch for pointing out improvements of prior results (wrt. [BFPZ22]).